# https://ku-cvlab.github.io/RAIN-GS

### v1

@@ 6.3 Qualitative and quantitative comparison
We compare our model with four methods: Plenoxels [39], InstantNGP-Base, InstantNGP-Big [22], and 3DGS [13]. To evaluate the effectiveness of our strategy when trained without SfM initialization, we make a thorough comparison with 3DGS trained using the original DSV random initialization. For DSV initialization, we uniformly sample points within three times the extent of the camera’s bounding box. We set the initial number of Gaussians to $N=1,000,000$. Our method achieves state-of-the-art results in all datasets, outperforming other models by a large margin.
6.3.1 Mip-NeRF360 dataset.
We conduct experiments on Mip-NeRF360 dataset and report our results in Table 1. Our method outperforms all the other methods when SfM point clouds are not available. Also, we provide qualitative results in Figure 7 on novel test views. Compared to InstantNGP and 3DGS trained with DSV random initialization, ours outperforms them in visual quality rendering fine-grained details with the least noise. From the comparison of (c) and (e), we can observe the effectiveness of our strategy, removing any unwanted high-frequency artifacts or floaters that are prevalent in the rendered depth shown in (c). More qualitative images are included in the supplementary material.
6.3.2 Tanks&Temples and Deep Blending dataset.
Table 2 shows the quantitative results on the Tanks&Temples and Deep Blending dataset which contain both indoor and outdoor scenes. Our method outperforms all other methods by a large margin, showing robustness regardless of domain shifts from indoor and outdoor scenes. Notably, in the “Train”, “Dr Johnson” and “Playroom” scenes, our method achieved a higher PSNR score than the ${}^{\dagger}$, which is trained with SfM initialization, showing the effectiveness of our method of relaxing accurately initialized point clouds for 3DGS training.
In Figure 7, we provide qualitative comparisons on the Tanks&Temples and Deep Blending datasets. Compared to 3DGS trained with DSV initialization, our method succeeds in rendering high-quality images and accurate depth. These results are also well-aligned with our motivation and methods explained in Section 4 and Section 5, reducing any unwanted high-frequency artifacts.
Table 3: Ablation on core components.
Table 4: Impact of $N$.
  Low-pass filter (LPF) init. Average
@@
Progressive(Convex) SLV 25.799 0.756 0.286
Progressive(Linear) SLV 25.898 0.754 0.290
Progressive(Concave) SLV 25.974 0.754 0.288
@@
[…]
w/o LPF w/ LPF
1M 22.190 23.937
100K 24.364 25.149
10K 24.798 25.959
1K 24.827 25.832
100 25.563 26.044
10 25.675 26.180
Table 4: Impact of $N$.
6.4 Ablation studies
6.4.1 Ablation on core components.
In Table 4, we validate the effectiveness of each component in our method. We compare our sparse-large-variance (SLV) initialization ($N=10$) against the original dense-small-variance (DSV) initialization ($N=1,000,000$). We also test our progressive Gaussian low-pass filter control against a constant filter ($s=0.3$) used in the original 3DGS [13]. To further validate our progressive Gaussian low-pass filter control formula ($s=HW/9\pi N$, ‘Ours’), we compare it against controlling the low-pass filter values sampled from linear, convex, and concave functions. Results verify our choice of dynamically adjusting the low-pass filter value based on the number of Gaussians, achieving the best performance. In addition, the results show that using both of our components: 1) sparse-large-variance (SLV) random initialization and 2) progressive Gaussian low-pass control is crucial showing the best performance. See Section A of the supplementary materials for more details of training with different low-pass filter value formulas.
6.4.2 Ablation on initial number of Gaussians.
In Table 4, we validate the effectiveness of our sparse-large-variance (SLV) initialization. Aligned with our analysis, the performance of both 3DGS trained with DSV initialization and our strategy increases as the initial number of Gaussians decreases. This further verifies our hypothesis that the sparse initialization of Gaussians leads to a more robust learning of the true distribution.
6.5 3DGS in sparse view settings
Due to the impressive performance shown by NeRF, various follow-ups have been proposed, specifically to succeed in learning NeRF with a limited number of images [23, 36, 12, 29, 16]. However, due to the requirement of accurate pose and initial point clouds of 3DGS, the task of successfully guiding 3D Gaussians with a limited number of images becomes a much more complicated task. In this section, we evaluate our strategy in training 3DGS with few-shot images, where SfM struggles to converge.
Figure 8: Qualitative comparison at sparse view setting.
We conduct our experiments on three scenes (bonsai, garden, kitchen) from the Mip-NeRF360 dataset, utilizing only 10% of the total training images. Quantitative and qualitative results of the sparse view setting are demonstrated in Table 5 and Figure 8. 3DGS trained with our strategy (Ours) outperforms 3DGS trained with DSV random initialization (3DGS (DSV)) both qualitatively and quantitatively, demonstrating the potential of our work to be extended to the task of training 3DGS with few-shot images. In Figure 8, it is shown that the rendered results of 3DGS trained with our strategy much more robustly models the real geometry of the scene.
Table 5: Few-shot Quantitative comparison on specific scene.
  Method bonsai garden kitchen
[…]
3DGS (DSV) 13.246 0.508 0.538 16.752 0.377 0.444 18.181 0.629 0.384
Ours 20.421 0.705 0.422 18.310 0.469 0.416 22.053 0.745 0.332
@@ 7 Conclusion
In this work, we introduce RAIN-GS, a novel optimization strategy that enables 3DGS to render high-quality images even from randomly initialized point clouds. By combining sparse-large-variance (SLV) random initialization and progressive Gaussian low-pass filter control, our strategy successfully guides 3D Gaussians to learn the low-frequency components first, which we demonstrate to be crucial for robust optimization. We evaluate the effectiveness of our strategy through comprehensive quantitative and qualitative comparisons, achieving state-of-the-art performance in all datasets. By effectively removing the strict reliance on accurate point clouds achieved from Structure-from-Motion (SfM), RAIN-GS opens up new possibilities for 3DGS in scenarios where accurate point clouds are not obtainable.
Appendix
@@
In the Appendix, we provide a more detailed analysis of our experiments and implementation details, together with additional results of rendered images using our strategy. Specifically, in Section 0.A, we provide the implementation details of how RAIN-GS (Relaxing Accurate INitialization Constraint for 3D Gaussian Splatting) can be applied to 3D Gaussian splatting (3DGS) [13]. In Section 0.B, we give a more detailed explanation of the convolution process with the Gaussian low-pass filter used in the rendering stage and how we choose the low-pass filter value $s$ (diagonal values of the covariance matrix $sI$ of the Gaussian low-pass filter) for our progressive Gaussian low-pass filter control. In Section 0.C, we explain the details of our toy experiment shown in Section 4.2: learning to regress 1D signal with multiple 1D Gaussians. In Section 0.D, we provide additional analysis including comparisons of computational resources with different methods and limitations of our strategy. In Section 0.E, we provide additional qualitative results. In Section 0.F, we show how our strategy can be applied to training 3DGS with a limited number of images.
Appendix 0.A Implementation Details
0.A.1 3DGS
We implement our strategy based on the official code of 3DGS [13]. As mentioned in Section 6.2, except for the hyperparameters for increasing spherical harmonic (SH) degrees and 3D Gaussian scale division ratio, we follow the original implementation and hyperparameters.
0.A.2 RAIN-GS
Our novel strategy consists of two components: 1) sparse-large-variance (SLV) random initialization and 2) progressive Gaussian low-pass filter control. In this section, we provide the pseudo codes for how we implement our strategy above 3DGS. We denote our implementation in the pseudo codes with a gray box.
0.A.2.1 Sparse-large-variance (SLV) random initialization.
@@ 0.A.2 RAIN-GS
Following the original implementation of 3DGS [13], we initialize random point clouds within the boundary defined as three times the size of the camera’s bounding box. Instead of the original dense-small-variance (DSV) random initialization where the initial number of Gaussians $N$ is set as $N>100K$, we only initialize 10 Gaussians as $N=10$. This simple change of code (expressed with a gray box in Algorithm 1) based on our analysis of sparse-large-variance (SLV) initialization, largely improves the performance of 3DGS. As the initial covariance of 3D Gaussian is defined based on the distances of the three nearest neighbors, sparse initialization leads to a larger initial covariance, resolving the requirement for additional modification.
Algorithm 1 Random point cloud initialization
@@
0.A.2.2 Progressive Gaussian low-pass filter control.
@@
Algorithm 2 Progressive Gaussian low-pass filter control
$M\leftarrow$ SfM Points $\triangleright$ Positions
$S,C,A\leftarrow$ InitAttributes() $\triangleright$ Covariances, Colors, Opacities
$i\leftarrow 0$triangleright Iteration Count
while not converged do
     $V,\hat{I},H,W\leftarrow$ SampleTrainingView() $\triangleright$ Camera $V$, Image, Height and Width
     if LowPassRefinementIteration(i) then
@@
     end if
     $I\leftarrow$ Rasterize($M$, $S$, $C$, $A$, $V$, $h$) $\triangleright$ Rasterization with low-pass filter
     $L\leftarrow Loss(I,\hat{I})$triangleright Loss
     $M$, $S$, $C$, $A$leftarrow Adam($\nabla L$) $\triangleright$ Backprop & Step
     if IsRefinementIteration($i$then
         for all Gaussians $(\mu,\Sigma,c,\alpha)$ in $(M,S,C,A)$ do
@@
                  RemoveGaussian()
              end if
              if $\nabla_{p}L>\tau_{p}$ then $\triangleright$ Densification
                  if $\|S\|>\tau_{S}$ then $\triangleright$ Over-reconstruction
                       SplitGaussian($\mu,\Sigma,c,\alpha$)
                  elsenormal-$\triangleright$ Under-reconstruction
                       CloneGaussian($\mu,\Sigma,c,\alpha$)
                  end if
              end if
         end for
     end if
     $i\leftarrow i+1$
end while
0.A.3 Ablation on core components details

### v2

@@ 6.4 Qualitative and quantitative comparison
In Table 4, we compare our model with Plenoxels [35], InstantNGP-Base, InstantNGP-Big [25], and 3DGS [1] on the Mip-NeRF360 dataset [21], Tanks&Temples dataset [44], and Deep Blending dataset [45]. All NeRF methods [35, 25, 20] are trained without SfM points. We evaluate 3DGS [1] and ours trained with SfM-initialized and randomly initialized point cloud. For ours trained with SfM-initialized point cloud, we evaluate utilizing different clustering methods of selecting the top 10% with the least reprojection error (denoted as <10%) or using cluster centers (denoted as cluster) obtained through HDBSCAN [46].
In Figure 4, we provide qualitative results on novel test views. We compare with InstantNGP-Base ((a)), 3DGS trained with random point cloud ((b)) and SfM initialized point cloud ((c)), and ours trained with random point cloud ((d)). By comparing (b) and (d), we demonstrate the effectiveness of our strategy, which removes any unwanted high-frequency artifacts or floaters and also successfully captures scene objects located far away from the viewpoint. By comparing (c) and (d), ours trained with random point cloud show better results than 3DGS trained with SfM initialized point cloud, capturing fine details more effectively and rendering the background more accurately. More qualitative images are included in Section E in the supplementary materials.
6.5 Ablation studies
Table 5: Ablation on core components
@@ 6.5 Ablation studies
Ours SLV 27.228 0.807 0.229
In Table 5, we validate the effectiveness of each component in our method trained in the Mip-NeRF360 dataset [21] with randomly initialized point cloud. The comparison of Sparse-large-variance (SLV) and dense-small-variance (DSV) initialization is done by using $N=10$ and $N=1,000,000$ respectively. We compare our progressive Gaussian low-pass filter control method with using a constant low-pass filter value ($s=0.3$) as done in the original 3DGS [1]. We also compare utilizing our novel ABE-Split method and the original split algorithm utilized in 3DGS. All three ablations show that using our components achieves significantly better results and using all three key components achieves the best results. More detailed and various ablation studies can be found in Section D.2 in the supplementary materials.
@@ 7 Conclusion
In this work, we introduce RAIN-GS (Relaxing Accurate INitialization Constraint for 3D Gaussian Splatting), a novel optimization strategy that effectively removes the constraint for accurate initialization. By combining sparse-large-variance (SLV) initialization, progressive Gaussian low-pass filter control, and the adaptive bound-expanding split (ABE-Split) algorithm, our strategy successfully guides 3D Gaussians to prioritize the learning of low-frequency components and encourages the Gaussians to transport to further locations from the initialized locations. RAIN-GS effectively removes the strict reliance on accurate point cloud achieved from Structure-from-Motion (SfM), opening up new possibilities for 3DGS in scenarios where achieving accurate point cloud is challenging.
Supplementary materials
@@
In the supplementary materials, we provide a more detailed analysis of our experiments and implementation details, together with additional results of rendered images using our strategy. Specifically, in Section A, we provide the implementation details of how our RAIN-GS (Relaxing Accurate INitialization Constraint for 3D Gaussian Splatting) can be applied to original 3D Gaussian splatting (3DGS) [1]. In Section B, we give a more detailed explanation of the convolution process with the Gaussian low-pass filter used in the rendering stage and how we choose the low-pass filter value $s$ (diagonal values of the covariance matrix $sI$ of the Gaussian low-pass filter) for our progressive Gaussian low-pass filter control. In Section C, we explain the details of our toy experiment shown in Section 5.1 of our main paper: learning to regress 1D signal with multiple 1D Gaussians. In Section D, we provide additional analysis including comparisons of computational resources with different methods and limitations of our strategy. In Section E, we provide additional qualitative results. In Section F, we show how our strategy can be applied to training 3DGS with a limited number of images.
Appendix A Implementation Details
A.1 RAIN-GS
We implement our strategy based on the official code of 3DGS [1]. As mentioned in Section  6.2 of our main paper, except for the hyperparameters for increasing spherical harmonic (SH) degrees and 3D Gaussian scale division ratio, we follow the original implementation details and hyperparameters of 3DGS [1]. We start increasing SH degree every 1,000 steps after 5,000 steps and we lower 3D Gaussian scale division ratio 1.6 to 1.4. Our novel strategy consists of three components: 1) sparse-large-variance (SLV) random initialization, 2) progressive Gaussian low-pass filter control and 3) ABE-Split. While 3DGS begins with the SFM point cloud which provides a coarse approximation of the scene, RAIN-GS starts with significantly fewer Gaussians and without any coarse information. Consequently to achieve an initial coarse approximation, it requires more time and steps to densify the Gaussians. Thus we add warm-up phase in training and our three main components are only applied during the warm-up phase. Specifically, the traditional 3DGS training can be divided into two main phases: (1) densification and (2) optimization. The densification phase typically involves about 15,000 steps, followed by 15,000 steps of optimization. The learning rate of the mean $\mu$ start with $1.6e^{-4}$ and exponentially decay to $1.6e^{-6}$ during densification phase. For RAIN-GS, however, the optimization phase is reduced to 5,000 steps, and 10,000 steps of warm-up phase is introduced prior to densification phase. While warm-up phase, the learning rate of the mean $\mu$ is kept $1.6e^{-4}$ and our componets are applied to help the Gaussians learn a coarse approximation. Then, the remaining process follows the same learning process as the original 3DGS.
A.1.1 Sparse-large-variance (SLV) random initialization
@@
Following the original implementation of 3DGS [1], we initialize random point cloud within the boundary defined as three times the size of the camera’s bounding box. Instead of the original dense-small-variance random initialization where the initial number of Gaussians $N$ is set as $N>100K$, we only initialize 10 Gaussians as $N=10$. As the initial covariance of 3D Gaussian is defined based on the distances of the three nearest neighbors, sparse initialization leads to a larger initial covariance, resolving the requirement for additional modification. This simple change of code (expressed with a gray box in Algorithm 1) largely improves the performance of 3DGS based on our analysis of sparse-large-variance initialization. An illustrations of SfM initialization, dense-small-variance (DSV) initialization and sparse-large-variance (SLV) initialization are shown in Figure 5.
Algorithm 1 Sparse-large-variance (SLV) random initialization
@@
A.1.2 Progressive Gaussian low-pass filter control
@@
Algorithm 2 Progressive Gaussian low-pass filter control & ABE-Split
1:$M\leftarrow$ SfM Points $\triangleright$ Positions
2:$S,C,A\leftarrow$ InitAttributes() $\triangleright$ Covariances, Colors, Opacities
3:$s\leftarrow$ Scene Extent
4:$B_{0}\leftarrow$ Center Point of Bounding Box
5:$i\leftarrow 0$triangleright Iteration Count
6:while not converged do
7:     $V,\hat{I},H,W\leftarrow$ SampleTrainingView() $\triangleright$ Camera $V$, Image, Height and Width
8:     if LowPassRefinementIteration(i) then
@@
10:     end if
11:     $I\leftarrow$ Rasterize($M$, $S$, $C$, $A$, $V$, $h$) $\triangleright$ Rasterization with low-pass filter
12:     $L\leftarrow Loss(I,\hat{I})$triangleright Loss
13:     $M$, $S$, $C$, $A$leftarrow Adam($\nabla L$) $\triangleright$ Backprop & Step
14:     if IsRefinementIteration($i$then
15:         for all Gaussians $(\mu,\Sigma,c,\alpha)$ in $(M,S,C,A)$ do
@@
17:                  RemoveGaussian()
18:              end if
19:              if $\nabla_{p}L>\tau_{p}$ then $\triangleright$ Densification
20:                  if $\|S\|>\tau_{S}$ then $\triangleright$ Over-reconstruction
21:                       SplitGaussian($\mu,\Sigma,c,\alpha$)
22:                       $\mu^{\prime}\leftarrow\lambda_{s}s\cdot(\mu-B_{0})+B_{0}$
23:                       CloneGaussian($\mu^{\prime},\Sigma,c,\alpha$) $\triangleright$ ABE-Split
24:                  $\triangleright$ Under-reconstruction
25:                       CloneGaussian($\mu,\Sigma,c,\alpha$)
26:                  end if
27:              end if
28:         end for
29:     end if
30:     $i\leftarrow i+1$
31:end while
A.1.3 ABE-Split
In addition to the split of 3DGS, the ABE-Split generates an additional Gaussian and transport it to a distant location. Specifically, new position is determined by calculating the vector between the center of bounding box and the original Gaussian position, then multiplying it by 0.3 times the scene extent. The pseudo code for this process is provided in line 21 and 22 of Algorithm 2. Due to our SLV initialization and progressive Gaussian low-pass filtering, the Gaussians are sparse and have large covariance during the early steps of training, resulting in a majority of the Gaussians only undergoing the split process. Thus, the ABE-Split can transport a sufficient number of Gaussians in early stage.
Appendix B Proof
B.1 Proof on radius of a Gaussian convolved with a low-pass filter
As mentioned in Section 5.2 of our main paper, the 3D Gaussians $G_{i}$ is projected to 2D Gaussians $G_{i}^{\prime}$ in the screen space as follows:
Following the convolution process, 3DGS estimates the projected 2D Gaussian’s area to identify its corresponding screen tiles. This is done by calculating $k$ times the square root of the larger eigenvalue of $(\Sigma_{i}^{\prime}+sI)$, which represents the radius of the approximated circle, and $k$ is the hyperparameter that determines the confidence interval of the 2D Gaussian. Figure 6 illustrates the low-pass filter’s effect, where the projected Gaussian is splatted to wider areas in (b) compared to (a).
B.2 Proof on progressive low-pass filter size
In Section 5.2 of our main paper, we define the value $s$ for our progressive Gaussian low-pass filter control based on the fact that the area of the projected 2D Gaussians is at least $9{\pi}s$. As the area of the projected 2D Gaussian is defined as the circle whose radius is $k$ times the square root of the larger eigenvalue of $(\Sigma_{i}^{\prime}+sI)$, we have to first calculate the eigenvalues of $(\Sigma_{i}^{\prime}+sI)$. If we define the eigenvalues of $\Sigma_{i}$ as $\lambda_{i1},\lambda_{i2}$, since the eigenvalue of $sI$ is $s$, the eigenvalues of $(\Sigma_{i}^{\prime}+sI)$ can be defined as $\lambda_{i1}+s,\lambda_{i2}+s$. This leads to the following proof:
[…]
Appendix C Toy experiments
C.1 Implementation details
In this section, we provide a detailed explanation of the 1D regression toy experiment shown in Section 5.1 of our main paper. We design the experiment to learn a target distribution $Y(x)$, where $x$ is in the range $[0,10,000)$, using $N$ 1D Gaussians. Each of the $N$ 1D Gaussians includes its mean $\mu_{i}$ and variance $\sigma_{i}^{2}$ as learnable parameters to model a random 1D signal $Y(x)$ as the ground truth distribution. The target distribution $Y(x)$ is generated from random 10 1D Gaussian distributions. By blending $N$ 1D Gaussians with a learnable weight $w_{i}$, we train the parameters $\{\mu_{i},\sigma_{i},w_{i}\}$ through the L1 loss between the blended distribution from $N$ 1D Gaussians and $Y(x)$ as follows:
Figure 7: Toy experiment to analyze different initialization methods. This figure visualizes the result of our toy experiment predicting the target distribution using a collection of 1D Gaussians, starting from different initialization methods. Dense-small-variance (DSV) and dense-large-variance (DLV) initialize 1,000 1D Gaussians, where DLV is initialized with large variances by adding $s$ to the initial variance. Small-large-variance (SLV) initializes 15 1D Gaussians with the same $s$ added to the initial variance. We can observe the over-fast convergence of high-frequency components on DSV initialization, which is resolved in the DLV initialization but fails to converge due to fluctuation. SLV initialization resolves both problems, learning the low-frequency components first and also converging to successfully model the target distribution.
C.2 Analysis on toy experiments
As mentioned in Section 5.1, we conduct a toy experiment in a simplified 1D regression task to examine how DSV and SLV initialization influences the optimization process. As shown in the top row of Figure 7, training with the DSV random initialization exhibits a tendency towards over-fast convergence on high frequencies. This is evident in the model’s ability to capture high-frequency components after only 10 steps. However, this rapid focus on high-frequencies leads to undesired high-frequency artifacts in the final reconstruction (zoomed-in box at 1,000 steps). We hypothesize that this behavior arises from the small initial variances of the dense Gaussians, which restrict each Gaussian to model a very localized region, increasing the susceptibility to overfitting.
To verify our hypothesis, we repeat the experiment with a value $s$ added to the initial variance changing our initialization to dense-large-variance (DLV). Note that this modification effectively encourages the Gaussians to learn from wider areas. The results (middle row of Figure 7) support our hypothesis: the learned signal at 10 steps demonstrates a prioritization of low-frequency components when compared to DSV initialization. However, even with the ability to prune Gaussians by learning weights as $w_{i}=0$, the dense initialization causes fluctuations in the learned signal, preventing convergence after 1,000 steps. These observations highlight a crucial point: while learning from wider areas (enabled by a large variance) is necessary to prioritize low-frequency components, a dense initialization can still lead to instability and convergence issues.
C.3 Analysis on dense-large-variance (DLV) initialization in 3DGS
We compare the influence of different dense initialization methods through the toy experiment of learning to regress a random 1D signal with multiple 1D Gaussians. Through the toy experiment, we analyzed that while Gaussian learning from wider areas (enabled by the large initial variance) leads to prioritizing low-frequency components, the dense initialization results in instability and convergence issues. As shown in Figure 7, starting from Gaussians with large variance (dense-large-variance (DLV)) leads to prioritizing the learning of low-frequency components, compared to dense-small-variance (DSV) initialization, observed in the coarse signal in the bottom row of step 10.
Table 6: Quantitative comparison on Mip-NeRF360 dataset with various initialization methods. We compare our method with the DSV random initialization method described in the original 3DGS [1], DLV random initialization, and SLV random initialization method. We report PSNR, SSIM, LPIPS.
Method Outdoor Scene
bicycle flowers garden stump treehill
[…]
Method Indoor Scene Average
room counter kitchen bonsai
[…]
Appendix D Additional analysis
D.1 Analysis on prioritized learning of low-frequency components
In Section 4 of our main paper, we find that SfM initialization guides 3D Gaussians with a coarse approximation of the true distribution to robustly learn the remaining high-frequency components. In Section 5.1 of our main paper, we then show that sparse-large-variance (SLV) random initialization successfully guides the Gaussians to prioritize the learning of low-frequency components. To verify that our strategy successfully guides 3DGS to prioritize the learning of low-frequency components, we provide additional analysis of our strategy through the rendered images achieved from different steps of training.
Figure 9: Images from Figure 8 analyzed in the frequency domain.
As mentioned in Section 4 of our main paper, 3DGS (top row) starts from a coarse approximation of the signal, which can be observed by the low-frequency components already being captured in step 100 ((b), top row). This is more evident when compared to the transformed signal of 3DGS (DSV) and Ours at step 100 ((b), middle and bottom row), where the low-frequency components are largely different from the GT signal. Aligned with our analysis of the toy experiment in Section 5.1 of our main paper, DSV random initialization exhibits a tendency towards over-fast convergence on high-frequencies. This can be observed by comparing (b) and (c), where the high-frequency components are quickly learned in (c). This leads to high-frequency artifacts after optimization, as shown in the rendered image (step 30,000, middle row) in Figure 8.
D.2 Ablation study on warm up phase
Ablation settings for low-pass filter control.